# algorithmic

### v1

@@ Appendix B AMP: closing on the marginals
The rBP iteration works with $\mathcal{O}(d^{2})$ truncated marginals on the edges of the factor graph, but can be approximated by an AMP iteration operating on the $d+n_{X}+n_{Y}$ full marginals of the nodes. This is possible since $J^{w,t}_{z,i\to ij},\;K^{w,t}_{z,i\to ij}$ and $J^{v,t}_{z,j\to ij},\;K^{v,t}_{z,j\to ij}$ depend only weakly on the target factor node. However one can not naively neglect the dependence of $\hat{v}^{z,t}_{j\to ij}$ and $\hat{w}^{z,t}_{i\to ij}$ on the target node, since a consistent expansion in $\mathcal{O}(n_{z}^{-\frac{1}{2}})$ results in the $\mathcal{O}(1)$ Onsager reaction terms which need to be taken into account in the estimators of the marginals’ means.
First, consider $J^{w,t}_{z,i},\;K^{w,t}_{z,i}$ and $J^{v,t}_{z,j},\;K^{v,t}_{z,j}$ which we get by not excluding the $k=j$ term (or $k=i$ respectively) from the summation in (A.12-A.11). For example,
$J^{w,t}_{z,i}$ $=\frac{\lambda_{z}}{\sqrt{n_{z}}}\sum_{k}^{d}S^{z}_{ik}\hat{v}^{z,t}_{k\to ik}.$ (B.1)
@@
$J^{w,t}_{z,i\to ij}-J^{w,t}_{z,i}$ $=-\frac{\lambda_{z}}{\sqrt{n_{z}}}S^{z}_{ij}\hat{v}^{z,t}_{j\to ij}=\mathcal{O}(n_{z}^{-\frac{1}{2}})$ (B.2)
$J^{v,t}_{z,j\to ij}-J^{v,t}_{z,j}$ $=-\frac{\lambda_{z}}{\sqrt{n_{z}}}S^{z}_{ij}\hat{w}^{z,t-1}_{i\to ij}=\mathcal{O}(n_{z}^{-\frac{1}{2}})$ (B.3)
and $\mathcal{O}(n^{-1}_{z})$ in the cases of $K^{w,t}_{z,i},K^{v,t}_{z,j}$, all negligible at $n_{z}\gg 1$. Note that here we also assumed that at each time-step, the $\hat{v}^{t}$ are updated first based on $\hat{w}^{t-1}$, and then the $\hat{w}^{t}$ based on $\hat{v}^{t}$, as discussed in Section 2.1. Next we want to replace also the means by target-independent versions $\hat{v}^{z,t}_{j}$ and $\hat{w}^{z,t}_{i}$ and so the variances by $\hat{\sigma}^{z,t}_{v,j}$ and $\hat{\sigma}^{z,t}_{w,j}$ The errors we make with this replacement,
$\hat{w}^{z,t}_{i\to ij}-\hat{w}^{z,t}_{i}$ $=f^{w}_{\mathrm{in}}(K^{w,t}_{z,i\to ij},J^{w,t}_{z,i\to ij})-f^{w}_{\mathrm{in}}(K^{w,t}_{z,i},J^{w,t}_{z,i})$ (B.4)
$=-\frac{\lambda_{z}}{\sqrt{n_{z}}}S^{z}_{ij}\hat{\sigma}^{z,t}_{w,i\to ij}\hat{v}^{z,t}_{j\to ij}+\mathcal{O}(\frac{1}{n_{z}})$ (B.5)
$\hat{v}^{z,t}_{j\to ij}-\hat{v}^{z,t}_{j}$ $=-\frac{\lambda_{z}}{\sqrt{n_{z}}}S^{z}_{ij}\hat{\sigma}^{z,t}_{v,j\to ij}\hat{w}^{z,t-1}_{i\to ij}+\mathcal{O}(\frac{1}{n_{z}}),$ (B.6)
are relevant since plugging into (B.1) we get errors $\sim\frac{1}{n_{z}}(S^{z}_{ik})^{2}$ which have non-vanishing mean of $\mathcal{O}(\frac{1}{n_{z}})$; thus replacing each of the $d$ terms of the sum in $J^{w,t}_{z,i}$, and $n_{z}$ terms of the sum in $J^{v,t}_{z,j}$, results in a compound error of $\mathcal{O}(\frac{d}{n_{z}})$ and $\mathcal{O}(1)$ respectively. Therefore, the Onsager correction terms (B.5) and (B.6) need to be added to the linear source terms $J^{w,t}_{z,i}$ and $J^{v,t}_{z,j}$ of the AMP iteration, yielding Equations (12) to (15) in the main text.
@@ B.1 Denoising functions for Gaussian priors
$\mathcal{W}(x,K,J)$ $=\frac{1}{2\pi\sqrt{\det\Sigma}}\exp\left(-\frac{1}{2}x^{T}\Sigma^{-1}x+J^{T}x-\frac{1}{2}x^{T}\mathrm{diag}(K)x\right)$ (B.7)
$=\frac{1}{2\pi\sqrt{\det\Sigma}}\exp\left(-\frac{1}{2}\left(x-\tilde{\Sigma}_{K}J\right)^{T}\tilde{\Sigma}_{K}^{-1}\left(x-\tilde{\Sigma}_{K}J\right)+\frac{1}{2}J^{T}\tilde{\Sigma}_{K}J\right)$ (B.8)
where $\tilde{\Sigma}_{K}=(\Sigma^{-1}+\mathrm{diag}(K))^{-1}$ To obtain $f^{v^{z}}_{\mathrm{in}}(K,J)=\partial_{J_{1/2}}\log\int\text{d}x\;\mathcal{W}^{v}(x,K,J)$ in the two-dimensional case, we use that the mean of a distribution is equal to the mean of its marginals, such that
$f^{v^{z}}_{\mathrm{in}}(K_{1},K_{2},J_{1},J_{2})$ $=(\tilde{\Sigma}_{K}J)_{z}=\tilde{\Sigma}_{zz}(K)J_{z}+\tilde{\Sigma}_{z\bar{z}}(K)J_{\bar{z}}.$ (B.9)
@@
$\left(\begin{matrix}\tilde{\Sigma}_{XX}(K)&\tilde{\Sigma}_{XY}(K)\\ \tilde{\Sigma}_{YX}(K)&\tilde{\Sigma}_{YY}(K)\end{matrix}\right)$ $=\frac{\det\Sigma}{(\sigma_{v^{X}}^{2}+K_{2}\det\Sigma)(\sigma_{v^{Y}}^{2}+K_{1}\det\Sigma)-c_{v}^{2}}\left(\begin{matrix}\sigma_{v^{X}}^{2}+K_{2}\det\Sigma&c_{v}\\ c_{v}&\sigma_{v^{Y}}^{2}+K_{1}\det\Sigma\end{matrix}\right),$ (B.10)
@@
$f^{w^{z}}_{\mathrm{in}}(K,J)=\frac{J}{K+\sigma_{w^{z}}^{-2}}.$ (B.11)
@@ B.2 Denoising function for Rademacher-Bernoulli prior
$=\log\left[\rho_{z}\cosh(J)e^{-\frac{1}{2}K}+(1-\rho_{z})\right]$ (B.12)
@@
$=\frac{\rho_{z}\sinh(J)e^{-\frac{1}{2}K}}{\rho_{z}\cosh(J)e^{-\frac{1}{2}K}+(1-\rho_{z})}.$ (B.13)
$=\frac{\rho_{z}\tanh(J)}{\rho_{z}+\frac{2(1-\rho_{z})}{\exp(J-\frac{1}{2}K)+\exp(-J-\frac{1}{2}K)}}$ (B.14)
where the last version (B.14) is stable against floating point overflows in numpy, that is it avoids any np.nan by avoiding 0*np.inf or 0/0 or np.inf/np.inf to occur, and used in the numerical implementations. For the derivative, a numerically benign version is
@@
$\quad\frac{\rho_{z}}{\left(\frac{1-\rho_{z}}{r(K,J)}+2\rho_{z}(1-\rho_{z})+\rho_{z}^{2}r(K,J)\right)}-\frac{\rho_{z}}{1-\rho_{z}+\rho_{z}r(K,J)}$ (B.15)
@@ B.3 Initialization of AMP
For sparse priors, AMP is known to have convergence problems for small noise at finite size, and when the trajectory leaves the proximity of the Nishimori line. Drift away from the Nishimori line arises in particular due to finite size noise close to the first-order transition. While also caused by additional factors such as nonzero mean of the data (Caltagirone et al., 2014) and there exist principled (Vila et al., 2015; Rangan et al., 2019) and non-principled (Sterk et al., 2023) mitigation techniques, these issues are importantly caused and partly avoidable by the initialization method.
@@
For small noise at finite size, that is $\sigma^{2}_{\xi}\sqrt{n}\sim\mathcal{O}(1)$, the expansion in $n^{-1/2}$ made in the derivation of rBP and AMP looses its accuracy. Here the well known spectral initialization is beneficial. It leaves the Nishimori line, but results in reliable convergence if the signal is strong (Celentano et al., 2023).
For moderate or larger noise, the average distance of the initialization from the Nishimori line can be minimized by rescaling a random sample from the prior such that the relation $Q_{w/v}^{z}=|M_{w/v}^{z}|$ holds on expectation for the given finite system size. This yields $\sigma^{2}_{\mathrm{init}}=\frac{\sigma^{2}_{\mathrm{prior}}}{n}$ for a vector in $\mathbb{R}^{n}$. Note that the distribution of the random overlap is still centered on zero, so this initialization can only minimize the average distance from the Nishimori line, not eliminate it, therefore we refer to it as "approximate Nishimori". To enforce the condition on the level of the single realization would require information about the ground-truth direction to enter the algorithm.
@@ Appendix C Linearized AMP: optimal spectral algorithm for weak recovery
     $\hat{v}\leftarrow\frac{\hat{v}}{||\hat{v}||}$
     $\hat{w}\leftarrow\frac{\hat{w}}{||\hat{w}||}$
@@
For priors of mean zero, we expand the rBP equations (A.14)-(A.17) around $\hat{w}^{z}_{i\to ij},\hat{v}^{z}_{j\to ij}=0$ to obtain a linearized rBP iteration, which is nothing but a power iteration of a linear operator. Again the the dimension of the operator can be reduced to $2d\times(n_{x}+n_{y})$ by the analogous steps as in Appendix B to obtain a power-iteration on the node level.
@@
$\left.\frac{\partial\hat{v}^{z,t}_{j\to ij}}{\partial\hat{w}^{z,t-1}_{k\to kj}}\right|_{w=0}$ $=\left.\frac{\partial f^{v^{z}}_{\mathrm{in}}}{\partial J^{z}}\right|_{w=0}\left.\frac{\partial J^{v,t-1}_{z,j\to ij}}{\partial\hat{w}^{z,t-1}_{k\to kj}}\right|_{w=0}=\sigma^{2}_{v^{z}}\frac{\lambda_{z}}{\sqrt{n_{z}}}S^{z}_{kj}\qquad\forall k\neq i$ (C.1)
$\left.\frac{\partial\hat{v}^{z,t}_{j\to ij}}{\partial\hat{w}^{\bar{z},t-1}_{k\to kj}}\right|_{w=0}$ $=\left.\frac{\partial f^{v^{z}}_{\mathrm{in}}}{\partial J^{\bar{z}}}\right|_{w=0}\left.\frac{\partial J^{v,t-1}_{\bar{z},j\to ij}}{\partial\hat{w}^{\bar{z},t-1}_{k\to kj}}\right|_{w=0}=c_{v}\frac{\lambda_{\bar{z}}}{\sqrt{n_{\bar{z}}}}S^{\bar{z}}_{kj}\qquad\forall k$ (C.2)
$\left.\frac{\partial\hat{w}^{z,t}_{i\to ij}}{\partial\hat{v}^{z,t}_{k\to ik}}\right|_{v=0}$ $=\left.\frac{\partial f^{w^{z}}_{\mathrm{in}}}{\partial J^{z}}\right|_{v=0}\left.\frac{\partial J^{w,t}_{z,i\to ij}}{\partial\hat{v}^{z,t}_{k\to ik}}\right|_{v=0}=\sigma^{2}_{w^{z}}\frac{\lambda_{z}}{\sqrt{n_{z}}}S^{z}_{ik}\qquad\forall k\neq j$ (C.3)
@@
$\hat{w}^{z,t}_{i}-\hat{w}^{z,t}_{i\to ij}$ $=\frac{\lambda_{z}^{2}}{n_{z}}\sigma^{2}_{v^{z}}\sigma^{2}_{w^{z}}\left(\sum_{k\neq j}S^{z}_{ik}S^{z}_{ik}\hat{w}^{z,t-1}_{i\to ij}+S^{z}_{ij}\sum_{k}S^{z}_{kj}\hat{w}^{z,t-1}_{k\to kj}\right)$ (C.4)
$=\frac{\lambda_{z}^{2}}{n_{z}}\sigma^{2}_{v^{z}}\sigma^{2}_{w^{z}}\left(\hat{w}^{z,t-1}_{i\to ij}\sum_{k}(S^{z}_{ik})^{2}+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})\right)$ (C.5)
$\hat{v}^{z,t}_{j}-\hat{v}^{z,t}_{j\to ij}$ $=\frac{\lambda_{z}^{2}}{n_{z}}\sigma^{2}_{v^{z}}\sigma^{2}_{w^{z}}\left(\hat{v}^{z,t-1}_{j\to ij}\sum_{k}(S^{z}_{kj})^{2}+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})\right)$ (C.6)

### v2

@@ Appendix B AMP: closing on the marginals
The rBP iteration works with $\mathcal{O}(d^{2})$ truncated marginals on the edges of the factor graph, but can be approximated by an AMP iteration operating on the $d+n_{X}+n_{Y}$ full marginals of the nodes. This is possible since $J^{w,t}_{z,i\to ij},\;K^{w,t}_{z,i\to ij}$ and $J^{v,t}_{z,j\to ij},\;K^{v,t}_{z,j\to ij}$ depend only weakly on the target factor node. However one can not naively neglect the dependence of $\hat{v}^{z,t}_{j\to ij}$ and $\hat{w}^{z,t}_{i\to ij}$ on the target node, since a consistent expansion in $\mathcal{O}(n_{z}^{-\frac{1}{2}})$ results in the $\mathcal{O}(1)$ Onsager reaction terms which need to be taken into account in the estimators of the marginals’ means. The validity of this correction for the more general MTP model has been established rigorously by [43].
First, consider $J^{w,t}_{z,i},\;K^{w,t}_{z,i}$ and $J^{v,t}_{z,j},\;K^{v,t}_{z,j}$ which we get by not excluding the $k=j$ term (or $k=i$ respectively) from the summation in (43-42). For example,
$J^{w,t}_{z,i}$ $=\frac{\lambda_{z}}{\sqrt{n_{z}}}\sum_{k}^{d}S^{z}_{ik}\hat{v}^{z,t}_{k\to ik}.$ (49)
@@
$J^{w,t}_{z,i\to ij}-J^{w,t}_{z,i}$ $=-\frac{\lambda_{z}}{\sqrt{n_{z}}}S^{z}_{ij}\hat{v}^{z,t}_{j\to ij}=\mathcal{O}(n_{z}^{-\frac{1}{2}})$ (50)
$J^{v,t}_{z,j\to ij}-J^{v,t}_{z,j}$ $=-\frac{\lambda_{z}}{\sqrt{n_{z}}}S^{z}_{ij}\hat{w}^{z,t-1}_{i\to ij}=\mathcal{O}(n_{z}^{-\frac{1}{2}})$ (51)
and $\mathcal{O}(n^{-1}_{z})$ in the cases of $K^{w,t}_{z,i},K^{v,t}_{z,j}$, all negligible at $n_{z}\gg 1$. Note that here we also assumed that at each time-step, the $\hat{v}^{t}$ are updated first based on $\hat{w}^{t-1}$, and then the $\hat{w}^{t}$ based on $\hat{v}^{t}$, as discussed in Section 2.1. Next we want to replace also the means by target-independent versions $\hat{v}^{z,t}_{j}$ and $\hat{w}^{z,t}_{i}$ and so the variances by $\hat{\sigma}^{z,t}_{v,j}$ and $\hat{\sigma}^{z,t}_{w,j}$ . The errors we make with this replacement,
$\hat{w}^{z,t}_{i\to ij}-\hat{w}^{z,t}_{i}$ $=f^{w}_{\mathrm{in}}(K^{w,t}_{z,i\to ij},J^{w,t}_{z,i\to ij})-f^{w}_{\mathrm{in}}(K^{w,t}_{z,i},J^{w,t}_{z,i})$ (52)
$=-\frac{\lambda_{z}}{\sqrt{n_{z}}}S^{z}_{ij}\hat{\sigma}^{z,t}_{w,i\to ij}\hat{v}^{z,t}_{j\to ij}+\mathcal{O}(\frac{1}{n_{z}})$ (53)
$\hat{v}^{z,t}_{j\to ij}-\hat{v}^{z,t}_{j}$ $=-\frac{\lambda_{z}}{\sqrt{n_{z}}}S^{z}_{ij}\hat{\sigma}^{z,t}_{v,j\to ij}\hat{w}^{z,t-1}_{i\to ij}+\mathcal{O}(\frac{1}{n_{z}}),$ (54)
are relevant since plugging into (49) we get errors $\sim\frac{1}{n_{z}}(S^{z}_{ik})^{2}$ which have non-vanishing mean of $\mathcal{O}(\frac{1}{n_{z}})$; thus replacing each of the $d$ terms of the sum in $J^{w,t}_{z,i}$, and $n_{z}$ terms of the sum in $J^{v,t}_{z,j}$, results in a compound error of $\mathcal{O}(\frac{d}{n_{z}})$ and $\mathcal{O}(1)$ respectively. Therefore, the Onsager correction terms (53) and (54) need to be added to the linear source terms $J^{w,t}_{z,i}$ and $J^{v,t}_{z,j}$ of the AMP iteration, yielding Equations (12) to (15) in the main text.
@@ B.1 Denoising functions for Gaussian priors
$\mathcal{W}(x,K,J)$ $=\frac{1}{2\pi\sqrt{\det\Sigma}}\exp\left(-\frac{1}{2}x^{T}\Sigma^{-1}x+J^{T}x-\frac{1}{2}x^{T}\mathrm{diag}(K)x\right)$ (55)
$=\frac{1}{2\pi\sqrt{\det\Sigma}}\exp\left(-\frac{1}{2}\left(x-\tilde{\Sigma}_{K}J\right)^{T}\tilde{\Sigma}_{K}^{-1}\left(x-\tilde{\Sigma}_{K}J\right)+\frac{1}{2}J^{T}\tilde{\Sigma}_{K}J\right)$ (56)
where $\tilde{\Sigma}_{K}=(\Sigma^{-1}+\mathrm{diag}(K))^{-1}$ . To obtain $f^{v^{z}}_{\mathrm{in}}(K,J)=\partial_{J_{1/2}}\log\int\text{d}x\;\mathcal{W}^{v}(x,K,J)$ in the two-dimensional case, we use that the mean of a distribution is equal to the mean of its marginals, such that
$f^{v^{z}}_{\mathrm{in}}(K_{1},K_{2},J_{1},J_{2})$ $=(\tilde{\Sigma}_{K}J)_{z}=\tilde{\Sigma}_{zz}(K)J_{z}+\tilde{\Sigma}_{z\bar{z}}(K)J_{\bar{z}}.$ (57)
@@
$\left(\begin{matrix}\tilde{\Sigma}_{XX}(K)&\tilde{\Sigma}_{XY}(K)\\ \tilde{\Sigma}_{YX}(K)&\tilde{\Sigma}_{YY}(K)\end{matrix}\right)$ $=\frac{\det\Sigma}{(\sigma_{v^{X}}^{2}+K_{2}\det\Sigma)(\sigma_{v^{Y}}^{2}+K_{1}\det\Sigma)-c_{v}^{2}}$
$\quad\times\left(\begin{matrix}\sigma_{v^{X}}^{2}+K_{2}\det\Sigma&c_{v}\\ c_{v}&\sigma_{v^{Y}}^{2}+K_{1}\det\Sigma\end{matrix}\right),$ (58)
@@
$f^{w^{z}}_{\mathrm{in}}(K,J)=\frac{J}{K+\sigma_{w^{z}}^{-2}}.$ (59)
@@ B.2 Denoising function for Rademacher-Bernoulli prior
$=\log\left[\rho_{z}\cosh(J)e^{-\frac{1}{2}K}+(1-\rho_{z})\right]$ (60)
@@
$=\frac{\rho_{z}\sinh(J)e^{-\frac{1}{2}K}}{\rho_{z}\cosh(J)e^{-\frac{1}{2}K}+(1-\rho_{z})}.$ (61)
$=\frac{\rho_{z}\tanh(J)}{\rho_{z}+\frac{2(1-\rho_{z})}{\exp(J-\frac{1}{2}K)+\exp(-J-\frac{1}{2}K)}}$ (62)
where the last version (62) is stable against floating point overflows in numpy, that is it avoids any np.nan by avoiding 0*np.inf or 0/0 or np.inf/np.inf to occur, and used in the numerical implementations. For the derivative, a numerically benign version is
@@
$\quad\frac{\rho_{z}}{\left(\frac{1-\rho_{z}}{r(K,J)}+2\rho_{z}(1-\rho_{z})+\rho_{z}^{2}r(K,J)\right)}-\frac{\rho_{z}}{1-\rho_{z}+\rho_{z}r(K,J)}$ (63)
@@ B.3 Initialization of AMP
For sparse priors, AMP is known to have convergence problems for small noise at finite size, and when the trajectory leaves the proximity of the Nishimori line. Drift away from the Nishimori line arises in particular due to finite size noise close to the first-order transition. While also caused by additional factors such as nonzero mean of the data [57] and there exist principled [58, 59] and non-principled [60] mitigation techniques, these issues are importantly caused and partly avoidable by the initialization method.
@@
For small noise at finite size, that is $\sigma^{2}_{\xi}\sqrt{n}\sim\mathcal{O}(1)$, the expansion in $n^{-1/2}$ made in the derivation of rBP and AMP looses its accuracy. Here the well known spectral initialization is beneficial. It leaves the Nishimori line, but results in reliable convergence if the signal is strong [47].
For moderate or larger noise, the average distance of the initialization from the Nishimori line can be minimized by rescaling a random sample from the prior such that the relation $Q_{w/v}^{z}=|M_{w/v}^{z}|$ holds on expectation for the given finite system size. This yields $\sigma^{2}_{\mathrm{init}}=\frac{\sigma^{2}_{\mathrm{prior}}}{n}$ for a vector in $\mathbb{R}^{n}$. Note that the distribution of the random overlap is still centered on zero, so this initialization can only minimize the average distance from the Nishimori line, not eliminate it, therefore we refer to it as “approximate Nishimori”. To enforce the condition on the level of the single realization would require information about the ground-truth direction to enter the algorithm.
@@ Appendix C Linearized AMP: optimal spectral algorithm for weak recovery
  $v^\hat{v}\leftarrow\frac$$\hat$$|\hat{v}$||}
  $w^\hat{w}\leftarrow\frac$$\hat$$|\hat{w}$||}
@@
For priors of mean zero, we expand the rBP equations (45)-(48) around $\hat{w}^{z}_{i\to ij},\hat{v}^{z}_{j\to ij}=0$ to obtain a linearized rBP iteration, which is nothing but a power iteration of a linear operator. Again the the dimension of the operator can be reduced to $2d\times(n_{x}+n_{y})$ by the analogous steps as in B to obtain a power-iteration on the node level.
@@
$\left.\frac{\partial\hat{v}^{z,t}_{j\to ij}}{\partial\hat{w}^{z,t-1}_{k\to kj}}\right|_{w=0}$ $=\left.\frac{\partial f^{v^{z}}_{\mathrm{in}}}{\partial J^{z}}\right|_{w=0}\left.\frac{\partial J^{v,t-1}_{z,j\to ij}}{\partial\hat{w}^{z,t-1}_{k\to kj}}\right|_{w=0}=\sigma^{2}_{v^{z}}\frac{\lambda_{z}}{\sqrt{n_{z}}}S^{z}_{kj}\qquad\forall k\neq i$ (64)
$\left.\frac{\partial\hat{v}^{z,t}_{j\to ij}}{\partial\hat{w}^{\bar{z},t-1}_{k\to kj}}\right|_{w=0}$ $=\left.\frac{\partial f^{v^{z}}_{\mathrm{in}}}{\partial J^{\bar{z}}}\right|_{w=0}\left.\frac{\partial J^{v,t-1}_{\bar{z},j\to ij}}{\partial\hat{w}^{\bar{z},t-1}_{k\to kj}}\right|_{w=0}=c_{v}\frac{\lambda_{\bar{z}}}{\sqrt{n_{\bar{z}}}}S^{\bar{z}}_{kj}\qquad\forall k$ (65)
$\left.\frac{\partial\hat{w}^{z,t}_{i\to ij}}{\partial\hat{v}^{z,t}_{k\to ik}}\right|_{v=0}$ $=\left.\frac{\partial f^{w^{z}}_{\mathrm{in}}}{\partial J^{z}}\right|_{v=0}\left.\frac{\partial J^{w,t}_{z,i\to ij}}{\partial\hat{v}^{z,t}_{k\to ik}}\right|_{v=0}=\sigma^{2}_{w^{z}}\frac{\lambda_{z}}{\sqrt{n_{z}}}S^{z}_{ik}\qquad\forall k\neq j$ (66)
@@
$\hat{w}^{z,t}_{i}-\hat{w}^{z,t}_{i\to ij}$ $=\frac{\lambda_{z}^{2}}{n_{z}}\sigma^{2}_{v^{z}}\sigma^{2}_{w^{z}}\left(\sum_{k\neq j}S^{z}_{ik}S^{z}_{ik}\hat{w}^{z,t-1}_{i\to ij}+S^{z}_{ij}\sum_{k}S^{z}_{kj}\hat{w}^{z,t-1}_{k\to kj}\right)$ (67)
$=\frac{\lambda_{z}^{2}}{n_{z}}\sigma^{2}_{v^{z}}\sigma^{2}_{w^{z}}\left(\hat{w}^{z,t-1}_{i\to ij}\sum_{k}(S^{z}_{ik})^{2}+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})\right)$ (68)
$\hat{v}^{z,t}_{j}-\hat{v}^{z,t}_{j\to ij}$ $=\frac{\lambda_{z}^{2}}{n_{z}}\sigma^{2}_{v^{z}}\sigma^{2}_{w^{z}}\left(\hat{v}^{z,t-1}_{j\to ij}\sum_{k}(S^{z}_{kj})^{2}+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})\right)$ (69)